OpenStudy (daenio):

12. An undercover news reporter wanted to know what exactly Fast Food hamburgers were made of. It was found that 95% of the burgers were made from beef, the rest was made from a Mystery Meat! a) What is the probability that at least 18 out of 20 of hamburgers selected randomly are made of beef? [4]

OpenStudy (daenio):

I tried #12. I'm not sure if it's right though. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

This is another binomial distribution problem. P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 18) = (20 C 18)*(0.95)^(18)*(1-0.95)^(20-18) P(X = 18) = (20 C 18)*(0.95)^(18)*(0.05)^(20-18) P(X = 18) = (190)*(0.95)^(18)*(0.05)^2 P(X = 18) = (190)*(0.397214318)*(0.0025) P(X = 18) = 0.18867680105 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 19) = (20 C 19)*(0.95)^(19)*(1-0.95)^(20-19) P(X = 19) = (20 C 19)*(0.95)^(19)*(0.05)^(20-19) P(X = 19) = (20)*(0.95)^(19)*(0.05)^1 P(X = 19) = (20)*(0.3773536025)*(0.05) P(X = 19) = 0.3773536025 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 20) = (20 C 20)*(0.95)^(20)*(1-0.95)^(20-20) P(X = 20) = (20 C 20)*(0.95)^(20)*(0.05)^(20-20) P(X = 20) = (1)*(0.95)^(20)*(0.05)^0 P(X = 20) = (1)*(0.3584859224085)*(1) P(X = 20) = 0.3584859224085 ===================================================================== P(X >= 18) = P(X = 18) + P(X = 19) + P(X = 20) P(X >= 18) = 0.18867680105 + 0.3773536025 + 0.3584859224085 P(X >= 18) = 0.9245163259585 So the probability that at least 18 out of 20 of hamburgers selected randomly are made of beef is 0.9245163259585

OpenStudy (daenio):

Oh, nice! I had the numbers I just didn't know what to do with them. Thank you!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (daenio):

Hey, I sent an attachment. Would it be possible if you checked two other questions out? Would you prefer if I opened the questions in separate convos?

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What did you get for the answer of #1?

OpenStudy (daenio):

I got binomial.

jimthompson5910 (jim_thompson5910):

for which part?

OpenStudy (daenio):

a)

jimthompson5910 (jim_thompson5910):

Is replacement allowed?

OpenStudy (daenio):

Is replacement another form of distribution?

jimthompson5910 (jim_thompson5910):

no I meant, are the cards replaced or put back when you draw them?

OpenStudy (daenio):

It isn't mentioned, so I would guess no.

jimthompson5910 (jim_thompson5910):

so then a hypergeometric distribution is used

jimthompson5910 (jim_thompson5910):

a binomial distribution is only used if the probability of each trial is the same

OpenStudy (daenio):

Oh, so the next one is binomial then.

jimthompson5910 (jim_thompson5910):

yes that's what I'd go with too

OpenStudy (daenio):

Actually, I think the answer is uniform.

OpenStudy (daenio):

This is a uniform probability distribution since there can only be one outcome and the probability of rolling a single die is the same.

jimthompson5910 (jim_thompson5910):

Ah very true, wasn't thinking that

OpenStudy (daenio):

If it were binomial, there would be two possible outcomes, success or failure.

jimthompson5910 (jim_thompson5910):

yes very true, I was thinking of rolling a 3 or something

OpenStudy (daenio):

Why do you think a) is hypergeometric? I'm a bit perplexed.

jimthompson5910 (jim_thompson5910):

Because I'm reading that the hypergeometric distribution is like the binomial distribution, but replacements aren't made (so the probabilities will change)

OpenStudy (daenio):

This is what I read---> However, while the binomial distribution relies on independent trials, the hypergeometric distribution relies on dependent trials. The probability of success changes with each successive trial.

jimthompson5910 (jim_thompson5910):

that's basically what I'm reading as well

jimthompson5910 (jim_thompson5910):

so the second draw depends on the first draw because the cards aren't replaced

OpenStudy (daenio):

Hey so, is the last one binomial since the coin is re-tossed?

jimthompson5910 (jim_thompson5910):

hmm don't see one about a coin flip

OpenStudy (daenio):

Hey, finding the probability of drawing a certain colour of candy out of a box of candies would be an example of a binomial. This is similar since the question is looking for how many fours are rolled.

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean by similar question

OpenStudy (daenio):

I mean the question is similar to the example. The example mentions that finding the probability of a CERTAIN colour of candy whereas in the question (1c.), it's about rolling a CERTAIN number, which in this case is 4.

jimthompson5910 (jim_thompson5910):

ah i gotcha

jimthompson5910 (jim_thompson5910):

well if you defined success as landing on that color, then you could argue that it is a binomial distribution

OpenStudy (daenio):

Alright.

OpenStudy (daenio):

Hey, did you open the attachment I sent? It includes a graph. Not sure how I could sent it otherwise on this thread, lol.

jimthompson5910 (jim_thompson5910):

I don't see a graph in it

OpenStudy (daenio):

Oh, sorry. I mean chart, lol.

jimthompson5910 (jim_thompson5910):

oh yes, i see it

jimthompson5910 (jim_thompson5910):

where did you need help with on this one?

OpenStudy (daenio):

Well, I tried doing #6 but I'm not sure if I was doing it right.

OpenStudy (daenio):

I did this at first. n= 2+ 4+6+8+10+12/ 6= 7 p= 0.12+0.31+0.04+0.18+0.22+0.13/ 6 = 0.17 E(x)= np = 7 (0.17) = 1.19

OpenStudy (daenio):

Then, I tried this. E(x)= 2(0.12) + 4 (0.31) + 6 (0.04) + 8(0.18) + 10(0.22) + 12(0.13) = 6.92

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The answer of 6.92 is the correct answer, so your second answer and work for it is correct.

OpenStudy (daenio):

Thanks I submitted my work for this question using my first attempt to my teacher earlier and she replied that "your numbers are correct, but what you did with them is not".

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I see, I'm glad you got some of it correct.

OpenStudy (daenio):

Thanks, but now I'm a bit skeptical since the solution that I for the first attempt was 1.19.

OpenStudy (daenio):

that I got*

jimthompson5910 (jim_thompson5910):

The solution is still 6.92 The numbers in your first attempt are correct, it's just how you used them is incorrect which gave you the wrong answer.

OpenStudy (daenio):

OOOOOH. Ok, lol.